OpenStudy (anonymous):

Calculate the mean, median, and mode of the following set of data. Round to the nearest tenth. 3, 15, 3, 3, 1, 11, 2, 8, 3, 7, 15

OpenStudy (anonymous):

you didnt put that in order

OpenStudy (anonymous):

Don't give out direct answer, you have to help him/her through the problem...smh

OpenStudy (anonymous):

Night-Watcher do you know the mean?

OpenStudy (anonymous):

The mean is all the data added together, then its divided by the amount of numbers in the set of data.

OpenStudy (anonymous):

THANKS! Also I have a few more math questions too please if you dont mind

OpenStudy (anonymous):

In this case its: The total amount of all the numbers in the set of data divided by 11

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Find the probability distribution for the following function: P(x) = mc055-1.jpgfor x = 3, 5, and 7

OpenStudy (anonymous):

That looks pretty confusing...:/

OpenStudy (anonymous):

Answer choices: A) P(3) = mc055-11.jpg, P(5) = mc055-12.jpg, P(7) = mc055-13.jpg B) P(3) = mc055-8.jpg, P(5) = mc055-9.jpg, P(7) = mc055-10.jpg C) P(3) = mc055-5.jpg, P(5) = mc055-6.jpg, P(7) = mc055-7.jpg D)P(3) = mc055-2.jpg, P(5) = mc055-3.jpg, P(7) = mc055-4.jpg

OpenStudy (anonymous):

tell me about it...its algebra 2

OpenStudy (anonymous):

Are there suppose to be pics with this?

OpenStudy (anonymous):

Because i see .jpg

OpenStudy (anonymous):

no I have no idea why thats there

OpenStudy (anonymous):

You should take a screen shot of it. Also i might not have enough time to help because i have to go at 1:00

OpenStudy (anonymous):

those arent even the answer choices on here so idk why their there. DANG IT!

OpenStudy (anonymous):

but for the mean, median, and mode one i just take 71 divide by 11?

OpenStudy (anonymous):

I described the mean above. The mode is the number in a set of data that occurs most often. And for the median, place the numbers you are given in value order and find the middle number

OpenStudy (anonymous):

gotcha! thanks for the help